OpenStudy (jpsmarinho):

|X^2 - 5x + 5 |<= 1 Someone knows how to solve that?

OpenStudy (anonymous):

|X^2 - 5x + 5 |<=| 1| |X^2 - 5x + 5 | -|1|<=0 |X^2 - 5x + 5 -1|<=0

jimthompson5910 (jim_thompson5910):

|x^2 - 5x + 5 | <= 1 -1 <= x^2 - 5x + 5 <= 1 x^2 - 5x + 5 >= -1 and x^2 - 5x + 5 <= 1 Solve each inequality for x to find your answers.

OpenStudy (jpsmarinho):

@myko it's incredible! I never had saw this. Thanks. Could you, please ,do me a favor? After that, what I should do?

OpenStudy (anonymous):

my answer is not totally correct, becouse |x|-|y|>=|x-y|. It would get to complecated later: Folow @jim_thompson5910 better

OpenStudy (jpsmarinho):

Aah, I understand! Thanks @jim_thompson5910 !